OpenStudy (anonymous):

One solution of the equation 2x2 − x − 10 = 0 is 5\2 . What is the other solution?

OpenStudy (anonymous):

using vietta's formulas we can say that -b/a=s1+s2 where s1=5/2, 1/2-5/2=s2,so s2=--4/2,s=-2

OpenStudy (anonymous):

A more simple method may be to use the quadratic formula. Do you know that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

quadratic formula involves a lot of stuff, since we know the one solution of the equation vietta is the easiest way.. vietta can be proved using the quadratic formula, just by adding the two possible solutions but that's a lot of trivial mathwork

OpenStudy (anonymous):

using the quadratic formula will increase your calculation a lot.

OpenStudy (anonymous):

Not arguing that vietta's formula would be easier, but does the original poster understand vietta's formula and what you put? The quadratic formula is very simple and is not a lot of calculations.

OpenStudy (anonymous):

all right. both of you can get corrected solution.

OpenStudy (anonymous):

@Chris9k yeahp that's right, but you have to do more stuff.. anyway both ways are correct, but the easiest one is always better ;)

OpenStudy (anonymous):

I do agree @infinity_ Whichever way you can make the least mistakes, but also understand what you are doing. :-)

OpenStudy (anonymous):

that's right..

OpenStudy (anonymous):

yes, if you can understand why the first way is right, it can't be better.